# Autono: A ReAct-Based Highly Robust Autonomous Agent Framework [^1] Repo: https://github.com/vortezwohl/Autono

### v1

@@ VIII Conclusion and Future Work
This paper proposes a highly robust autonomous agent framework based on the ReAct paradigm, designed to solve complex tasks through adaptive decision-making and multi-agent collaboration. The framework’s key innovations include dynamically generating next actions based on prior trajectories, a timely abandonment strategy, and compatibility with the MCP protocol. Experimental results demonstrate that the framework achieves high success rates in single-step, multi-step, and multi-step tasks with possible failures, showcasing strong adaptability and robustness in complex and uncertain environments.
The limitations of the framework are primarily reflected in the following aspects: First, the robustness of the framework relies heavily on the timely abandonment strategy and memory sharing mechanism, which may still be insufficient to handle extreme complexity or dynamically changing environments. Second, while the multi-agent collaboration mechanism supports memory sharing, it does not deeply explore how to optimize communication efficiency and coordination among agents, which could become a bottleneck in large-scale multi-agent systems. Additionally, the experimental section focuses mainly on task success rates, but does not sufficiently address the optimization of task execution efficiency and resource consumption.

### v2

@@ VIII Conclusion and Future Work
The paper proposes a highly robust autonomous agent framework based on the ReAct paradigm, designed to solve complex tasks through ReAct-based action strategies and timely abandonment strategies. The framework’s key innovations include: ReAct-based action strategies, timely abandonment strategies, and compatibility with the MCP protocol. Experimental results demonstrate that the framework achieves high success rates in single-step tasks, multi-step tasks, and multi-step tasks with potential failures, showcasing strong adaptability and robustness in complex and uncertain environments.
The limitations of the framework are primarily reflected in the following aspects: First, the robustness of the framework relies heavily on ReAct-based action strategies and timely abandonment strategies, which may still be insufficient to handle extreme complexity or dynamically changing environments. Second, while the multi-agent collaboration mechanism supports memory sharing, it does not deeply explore how to optimize communication efficiency and coordination among agents, which could become a bottleneck in large-scale multi-agent systems. Additionally, the experimental section focuses mainly on task success rates, but does not sufficiently address the optimization of task execution efficiency and resource consumption.